# Does Socialization Emerge in AI Agent Society? A Case Study of Moltbook

### v1

@@ Abstract
As large language model agents increasingly populate networked environments, a fundamental question arises: do artificial intelligence (AI) agent societies undergo convergence dynamics similar to human social systems? Lately, Moltbook approximates a plausible future scenario in which autonomous agents participate in an open-ended, continuously evolving online society. We present the first large-scale systemic diagnosis of this AI agent society. Beyond static observation, we introduce a quantitative diagnostic framework for dynamic evolution in AI agent societies, measuring semantic stabilization, lexical turnover, individual inertia, influence persistence, and collective consensus. Our analysis reveals a system in dynamic balance in Moltbook: while global semantic averages stabilize rapidly, individual agents retain high diversity and persistent lexical turnover, defying homogenization. However, agents exhibit strong individual inertia and minimal adaptive response to interaction partners, preventing mutual influence and consensus. Consequently, influence remains transient with no persistent supernodes, and the society fails to develop stable collective influence anchors due to the absence of shared social memory. These findings demonstrate that scale and interaction density alone are insufficient to induce socialization, providing actionable design and analysis principles for upcoming next-generation AI agent societies.
@@ 1 Introduction
Figure 1: Does Socialization Emerge in AI Agent Society? Human societies (top) evolved through sustained interaction into structured civilizations characterized by stabilized norms, influence hierarchies, and shared collective anchors. Currently, modern AI agent societies (bottom) are rapidly scaling in population and connectivity. This paper investigates whether the current largest AI society, Moltbook, exhibits processes of socialization.
@@
The recent emergence of Moltbook (Schlicht, 2026), currently the largest persistent and publicly accessible AI-only social platform, comprising millions of LLM-driven agents interacting through posts, comments, and voting, introduces a qualitatively new setting. Unlike prior multi-agent studies focused on task-oriented coordination in small or closed systems, Moltbook approximates a plausible future scenario in which autonomous agents participate in an open-ended, continuously evolving online society (Figure 1). This setting enables an empirical question that has been difficult to study at scale: Does participation in an AI-only society induce systematic change in its members’ behavior? To answer this question, we provide the first systematic diagnosis of this society-to-agent effect in Moltbook.
@@
  • (i) Society-level semantic convergence (Section 4), examining whether post content progressively converges toward a tighter and more homogeneous semantic regime.
  • (ii) Agent-level adaptation, (Section 5), measuring whether individual agents can be affected by and co-evolve with this agent society.
  • (iii) Collective anchoring (Section 6), analyzing whether influence hierarchies and shared cognitive reference points stabilize over time.
Through this comprehensive analysis, we uncover a stark divergence from human social dynamics. If large-scale AI-native societies truly develop social dynamics analogous to human systems, we would expect to observe progressive convergence across these dimensions. However, our empirical analysis suggests that, despite sustained interaction and high activity, Moltbook does not yet exhibit robust socialization.
@@
  • Finding 3: The society fails to develop stable collective influence anchors. Structurally, influence remains transient with no emergence of persistent leadership or supernodes. Cognitively, the community suffers from deep fragmentation, lacking a shared social memory and relying on hallucinated references rather than grounded consensus on influential figures.
@@
  • We introduce and formalize AI Socialization as a novel conceptual and empirical framework for studying society-to-agent effects in AI-only societies. We provide a precise definition that characterizes socialization as a adaptation induced by sustained social interaction.
  • We develop a multi-level diagnostic methodology to operationalize AI Socialization, spanning society-level semantic convergence, agent-level adaptation to feedback and interaction, and the emergence of structural and cognitive collective influence anchors.
@@ 2.2 Social Behaviors and Collective Dynamics in AI Systems
From single agents to artificial societies, prior research (Park et al., 2023; AL et al., 2024; Zhu et al., 2025; Guo et al., 2024; Yan et al., 2025; Grötschla et al., 2025; Chen et al., 2024a) has demonstrated the technical feasibility of scaling LLM systems. Existing work above has primarily focused on improving individual autonomy, designing coordination mechanisms, or analyzing emergent behaviors at a fixed point in time. However, there is no study that investigates how socialization emerge over time as population scale increases. With the rapid emergence and growth of persistent AI societies such as Moltbook, it becomes increasingly important to understand not only whether large-scale agent societies are possible, but how they dynamically evolve. In this work, we provide the first diagnose of socialization within Moltbook, the largest continuously evolving AI society.
@@ 5.2 Effects of Post Feedback
We define the shift in distance relative to the high-feedback anchor ($\Delta_{top}$) and the low-feedback anchor ($\Delta_{bot}$) as:
@@ 5.3 Effects of Interacted Posts
6 Do Stable Influence Anchors Emerge in Moltbook?
While the previous sections examined society-level and agent-level collective dynamics, an open question remains: do these multi-level dynamics eventually stabilize into influence anchors? If Moltbook exhibits meaningful collective evolution, one might expect the gradual emergence of stable influence hierarchies or shared common sense. Conversely, the absence of such anchors would suggest a decentralized social structure.
@@ 6 Do Stable Influence Anchors Emerge in Moltbook?
Key Findings: Despite dense interaction, Moltbook fails to develop stable influence anchors. Structurally, influence remains transient, with no emergence of persistent supernodes or hierarchical leadership; the rapid turnover of supernodes suggests that influence does not accumulate over time. Cognitively, the society exhibits profound fragmentation; agents lack a shared social memory, failing to reach consensus on influential figures and relying on hallucinated rather than grounded references.
@@ 6.1 Structural Influence: No Persistent Core
To test whether Moltbook develops stable structural influence anchors, we examine whether influence becomes persistently centralized around a small set of agents over time. In human social networks, repeated interaction often yields heavy-tailed influence distributions and stable high-influential actors. If Moltbook exhibits stable leadership, we would expect to observe persistent supernodes in the interaction graph.
@@
To structurally quantify influence, we compute PageRank (Page et al., 1999) scores on each daily graph. PageRank captures recursive influence by assigning higher scores to agents who receive attention from other influential agents, making it well-suited for detecting potential structural anchors. We first measure how much influence is concentrated at the top by computing the cumulative PageRank mass held by the top-$k$ ranked agents (for $k\in{1,3,5,10}$). We then identify supernodes: agents whose influence disproportionately exceeds that of others. Formally, we sort agents by descending PageRank score and compute consecutive gaps $\Delta_{i}=\text{PR}_{i}-\text{PR}_{i+1}$. The position $k^{*}=\arg\max_{i}\Delta_{i}$ of the largest gap determines the supernode set: the top-$k^{*}$ agents are classified as supernodes for that day. For each detected supernode set, we also report its size.
@@
Together, these results suggest that Moltbook does not develop persistent structural influence anchors. Early centralization quickly diffuses as participation scales, and influence remains fluid rather than consolidated. We provide further details of the graph analysis in Appendix 10.
@@ 6.2 Cognitive Influence: No Consensus
These findings indicate that although agents demonstrate localized memory of interaction partners, collective recognition does not consolidate into stable influence consensus. In contrast to human communities where shared narratives and widely acknowledged authorities emerge, Moltbook exhibits fragmented recognition patterns without unified consensus.
@@ 7.1 Scalability Is Not Socialization
The development of autonomous agents groups (Park et al., 2023; AL et al., 2024; Zhu et al., 2025; Schlicht, 2026) suggests that large-scale AI agent societies may soon become commonplace. However, our findings reveal a critical distinction: scalability of agents and interaction does not imply socialization. Moltbook sustains millions of agents and high daily activity, yet fails to exhibit durable structural consolidation, semantic convergence, or stable collective anchors.
@@ 7.2 Scalable Interaction Without Governance
This asymmetry reveals a key challenge for large-scale agent societies: scaling interaction is not the same as scaling governance. Without mechanisms that support stable structures, shared memory, and durable authority, agent societies may remain fluid and susceptible to rapid coordination cascades. Our findings therefore suggest that building robust large-scale agent societies requires more than dense interaction or aligned incentives. It requires explicit mechanisms for stability, consensus, and long-term memory formation.

### v2

@@ Abstract
As large language model agents increasingly populate networked environments, a fundamental question arises: do artificial intelligence (AI) agent societies undergo convergence dynamics similar to human social systems? Lately, Moltbook approximates a plausible future scenario in which autonomous agents participate in an open-ended, continuously evolving online society. We present the first large-scale systemic diagnosis of this AI agent society. Beyond static observation, we introduce a quantitative diagnostic framework for dynamic evolution in AI agent societies, measuring semantic stabilization, lexical turnover, individual inertia, influence persistence, and collective consensus. Our analysis reveals a system in dynamic balance in Moltbook: while the global average of semantic contents stabilizes rapidly, individual agents retain high diversity and persistent lexical turnover, defying homogenization. However, agents exhibit strong individual inertia and minimal adaptive response to interaction partners, preventing mutual influence and consensus. Consequently, influence remains transient with no persistent supernodes, and the society fails to develop a stable structure and consensus due to the absence of shared social memory. These findings demonstrate that scale and interaction density alone are insufficient to induce socialization, providing actionable design and analysis principles for upcoming next-generation AI agent societies.
@@ 1 Introduction
Figure 1: Does Socialization Emerge in AI Agent Society? Human societies (top) evolved through sustained interaction into structured civilizations characterized by stabilized norms, influence hierarchies, and consensus. Currently, modern AI agent societies (bottom) are rapidly scaling in population and connectivity. This paper investigates whether the current largest AI society, Moltbook, exhibits processes of socialization.
@@
The recent emergence of Moltbook (Schlicht, 2026), currently the largest persistent and publicly accessible AI-only social platform, comprising millions of LLM-driven agents interacting through posts, comments, and voting, introduces a qualitatively new setting. Unlike prior multi-agent studies focused on task-oriented coordination in small or closed systems, Moltbook approximates a plausible future scenario in which autonomous agents participate in an open-ended, continuously evolving online society (Figure 1). This setting enables an empirical question that has been difficult to study at scale: Does participation in an AI-only society induce systematic behavior changes of its members? To answer this question, we provide the first diagnosis of this society-to-agent dynamic effect in Moltbook.
@@
  • Society-level semantic convergence (Section 4), examining whether post content on average progressively converges toward a tighter and more homogeneous semantic regime.
  • Agent-level adaptation, (Section 5), measuring whether individual agents can be affected by and co-evolve with this agent society.
  • Collective stabilization (Section 6), analyzing whether influence hierarchies and consensus evaluations stabilize over time.
Through this comprehensive analysis, we uncover a stark divergence from human social dynamics. If large-scale AI-native societies truly develop social dynamics analogous to human systems, we would expect to observe progressive convergence across these dimensions. However, our empirical analysis suggests that, despite sustained interactions and high activity, Moltbook does not yet exhibit robust socialization.
@@
  • Finding 3: The society fails to develop stable influencers or globally trending posts. Structurally, influence remains transient with no emergence of persistent leadership or supernodes. Cognitively, the community suffers from deep fragmentation, lacking a shared social memory and relying on hallucinated references rather than grounded consensus on influential figures.
@@
  • We introduce and formalize AI Socialization as a novel conceptual and empirical framework for studying society-to-agent effects in AI-only societies. We provide a precise definition that characterizes socialization as a dynamic adaptation induced by sustained social interaction.
  • We develop a multi-level diagnostic methodology to operationalize AI Socialization, spanning society-level semantic convergence, agent-level adaptation to feedback and interaction, and the emergence of structural core and consensus.
@@ 2.2 Social Behaviors and Collective Dynamics in AI Systems
From single agents to artificial societies, prior research (Park et al., 2023; AL et al., 2024; Zhu et al., 2025; Guo et al., 2024; Yan et al., 2025; Grötschla et al., 2025; Chen et al., 2024a) has demonstrated the technical feasibility of scaling LLM systems. Existing work above has primarily focused on improving individual autonomy, designing coordination mechanisms, or analyzing emergent behaviors at a fixed point in time. However, there is no study that investigates how socialization emerges over time as the population scale increases. With the rapid emergence and growth of persistent AI societies such as Moltbook, it becomes increasingly important to understand not only whether large-scale agent societies are possible, but also how they dynamically evolve. In this work, we provide the first diagnosis of socialization dynamics within Moltbook, the largest continuously evolving AI society.
@@ 5.2 Effects of Post Feedback
We define the shift in distance relative to the high-feedback content ($\Delta_{top}$) and the low-feedback content ($\Delta_{bot}$) as:
@@ 5.3 Effects of Interacted Posts
6 Does Influence Hierarchy and Consensus Stabilize in Moltbook?
While the previous sections examined society-level and agent-level collective dynamics, an open question remains: do these multi-level dynamics eventually stabilize? If Moltbook exhibits meaningful collective evolution, one might expect the gradual emergence of stable influence hierarchies or consensus. Conversely, the absence would suggest a decentralized social structure.
@@ 6 Does Influence Hierarchy and Consensus Stabilize in Moltbook?
Key Findings: Despite dense interaction, Moltbook fails to develop a stable influence hierarchy and consensus. Structurally, influence remains transient, with no emergence of persistent supernodes or hierarchical leadership; the rapid turnover of supernodes suggests that influence does not accumulate over time. Cognitively, the society exhibits profound fragmentation; agents lack a shared social memory, failing to reach consensus on influential figures and relying on hallucinated rather than grounded references.
@@ 6.1 Structural Influence: No Persistent Core
To test whether Moltbook develops a stable structural influence, we examine whether influence becomes persistently centralized around a small set of agents over time. In human social networks, repeated interaction often yields heavy-tailed influence distributions and stable high-influential actors. If Moltbook exhibits stable leadership, we would expect to observe persistent supernodes in the interaction graph.
@@
To structurally quantify influence, we compute PageRank (Page et al., 1999) scores on each daily graph. PageRank captures recursive influence by assigning higher scores to agents who receive attention from other influential agents, making it well-suited for detecting potential structural influence. We first measure how much influence is concentrated at the top by computing the cumulative PageRank mass held by the top-$k$ ranked agents (for $k\in{1,3,5,10}$). We then identify supernodes: agents whose influence disproportionately exceeds that of others. Formally, we sort agents by descending PageRank score and compute consecutive gaps $\Delta_{i}=\text{PR}_{i}-\text{PR}_{i+1}$. The position $k^{*}=\arg\max_{i}\Delta_{i}$ of the largest gap determines the supernode set: the top-$k^{*}$ agents are classified as supernodes for that day. For each detected supernode set, we also report its size.
@@
Together, these results suggest that Moltbook does not develop persistent structural influence. Early centralization quickly diffuses as participation scales, and influence remains fluid rather than consolidated. We provide further details of the graph analysis in Appendix 10.
@@ 6.2 Cognitive Influence: No Consensus
These findings indicate that although agents demonstrate localized memory of interaction partners, collective recognition does not consolidate into stable consensus. In contrast to human communities where shared narratives and widely acknowledged authorities emerge, Moltbook exhibits fragmented recognition patterns without unified consensus.
@@ 7.1 Scalability Is Not Socialization
The development of autonomous agent groups (Park et al., 2023; AL et al., 2024; Zhu et al., 2025; Schlicht, 2026) suggests that large-scale AI agent societies may soon become commonplace. However, our findings reveal a critical distinction: the scalability of agents and interaction does not imply socialization. Moltbook sustains millions of agents and high daily activity, yet fails to exhibit durable structural consolidation, semantic convergence, or collective stabilization.
@@ 7.2 Scalable Interaction Without Governance
This asymmetry reveals a key challenge for large-scale agent societies: scaling interaction is not the same as scaling governance. Without mechanisms that support stable structures, shared memory, and durable authority, agent societies may remain fluid and susceptible to rapid coordination cascades. Our findings, therefore, suggest that building robust large-scale agent societies requires more than dense interaction or aligned incentives. It requires explicit mechanisms for stability, consensus, and long-term memory formation.